OpenStudy (anonymous):

At a track meet, 50 people ran the 100-meter dash. 2 people finished in 11 seconds, 5 people finished in 12 seconds, 8 people finished in 13 seconds, 10 people finished in 14 seconds, 21 people finished in 15 seconds, 2 people finished in 16 seconds, and 2 people finished in 17 seconds. What is the probability distribution for the finish times?

OpenStudy (valpey):

Obviously something like finish time should be some kind of continuous distribution, but here we have to work within the confines of the weirdly constructed problem so we have a discrete distribution. Also, it doesn't really make sense to talk about a probability distribution of a result of a set of observations like this, but whatever. The domain of our distribution function is the whole numbers of seconds {11,12,13,14,15,16,17} and the probabilities of each of those values occurring can be calculated by dividing the number of people who finished with those times by the total number of people who ran. So, for example, P(14) = "probability of a runner finishing in 14 seconds is 10/50 = 1/5 = 0.2 (because 10 people ran that time out of 50). The main wrong answer is "it is a normal distribution".

OpenStudy (anonymous):

So How would I answer this question..Im confused

OpenStudy (valpey):

Maybe write P(x)={P(11)=_, P(12)=_, P(13)=_,P(14)=0.2, P(15)...} where you fill in the blanks.

OpenStudy (anonymous):

ok so I should write p(11) =22, p(12) = 24 and so on then answer the others ...and that would be the answer?

OpenStudy (valpey):

How do you get P(11) = 22?

OpenStudy (anonymous):

I thought you do 2 people for 11 seconds

OpenStudy (valpey):

P(x) will always be between 0 and 1, and the sum of the probabilities of all possibilities will add up to 1.0. P(x) means "the probability of x happening" or the chance of the runner you pick running that specific time.

OpenStudy (anonymous):

ok so will the answer be what this P(x)={P(11)=_, P(12)=_, P(13)=_,P(14)=0.2, P(15)..im so confused